OpenStudy (anonymous):

Which of the following equations has only one solution? x^2+ 4x + 4 = 0 x^2+ x = 0 x^2- 1 = 0

OpenStudy (lgbasallote):

do you mean one UNIQUE solution?

OpenStudy (lgbasallote):

because if it's just "one solution" then it should be none of the above because: linear equation = one solution quadratic equation = two solutiions cubic equation = three solutions and so on... and all of these are quadratic equations...

OpenStudy (lgbasallote):

anyway....factor out the equations and tell me what you got

OpenStudy (anonymous):

fjx^2+ 4x + 4 = 0

OpenStudy (anonymous):

x^2+ 4x + 4 = 0

OpenStudy (lgbasallote):

so factor that..

OpenStudy (anonymous):

x^2+4x+4=0 (x+2)^2=0 =-2?

OpenStudy (anonymous):

x^2+ x = 0 x^2+x=0 x=0? -1?

OpenStudy (anonymous):

x^2- 1 = 0 x^2-1=0 x^2=1 x=-1? 1?

OpenStudy (lgbasallote):

all are correct so which one has 1 solution?

OpenStudy (anonymous):

x^2+4x+4=0 :)

OpenStudy (lgbasallote):

right